Get the GCD and the LCM of 15np, 6n², and 39n²p
Answer by Edwin McCravy(20065)

Get the GCD and the LCM of 15np, 6n², and 39n²p
Break each down into their prime factors with no exponents:

15np = 3·5·n·p
6n² = 2·3·n·n
39n²p=3·13·n·n·p

Line those factors up so that all factors in each vertical column 
are the same, and draw a line underneath

15np  =   3·5·   n·  p
  6n² = 2·3·     n·n
39n²p =   3·  13·n·n·p

To get the LCM, bring every factor down under the bottom line, like this:

15np  =   3·5·   n·  p
  6n² = 2·3·     n·n
39n²p =   3·  13·n·n·p
  LCM = 2·3·5·13·n·n·p = 390n²p

To get the GCD, bring only the factors down under the bottom line,
if they appear on every line, like this:

15np  =   3·5·   n·  p
  6n² = 2·3·     n·n
39n²p =   3·  13·n·n·p
  GCD =   3     ·n     = 3n

Edwin